If f(x) = |3x| - 10 and g(x) = 2x^3, what is (g*f)( - 2)?
Answer by math_tutor2020(3821)


Answer: 64

Scratch Work
f(x) = |3x| - 10
f(-2) = |3*(-2)| - 10
f(-2) = |-6| - 10
f(-2) = 6 - 10
f(-2) = -4

g(x) = 2x^3
g(-2) = 2(-2)^3
g(-2) = 2(-8)
g(-2) = -16

(g*f)(-2) = g(-2)*f(-2)
(g*f)(-2) = -16*(-4)
(g*f)(-2) = 64
You can use Desmos or GeoGebra (among many other tools) as a way to confirm.